OpenStudy (anonymous):

MEDAL TO WHOEVER ANSWERS FIRST. GIven triangle ABC is congruent to triangle PQR, and m

OpenStudy (lukecrayonz):

3v+4=8v-6

OpenStudy (lukecrayonz):

Solve for v. input back into the original equations.

OpenStudy (anonymous):

i have no clue how to do that i am absolutely horrible at geometry/ algebra.

OpenStudy (lukecrayonz):

Watch KhanAcademy videos because that is basic 6th-7th grade math.

OpenStudy (anonymous):

you cant just give me the answer? it would make this a heck of a lot easier.

OpenStudy (lukecrayonz):

No I cannot, that's against OS rules.

OpenStudy (lukecrayonz):

3v+4=8v-6 You have to get v by itself. Do you know your first step

OpenStudy (anonymous):

subtract 3 from both sides?

OpenStudy (lukecrayonz):

What makes you say that

OpenStudy (anonymous):

i remember doing something like that

OpenStudy (mathstudent55):

3v + 4 = 8v - 6 You want all terms with v on the left side, and all terms with no variable on the right side. First, subtract 8v from both sides: 3v + 4 = 8v - 6 -8v -8v ----------------- -5v + 4 = -6 Now we need the 4 moved to the right side, so let's subtract 4 from both sides: -5v + 4 = -6 - 4 -4 --------------- -5v = -10 Since v is being multiplied by -5, we need to divide both sides by -5: -5v/(-5) = -10/(-5) v = 2 Now that we have v = 2, we insert v in the expressions used in the measure of the angles: m<B = 3v + 4 m<Q = 8v - 6 m<B = 3(2) + 4 m<Q = 8(2) - 6 m<B = 6 + 4 m<Q = 16 - 6 m<B = 10 m<Q = 10

OpenStudy (anonymous):

omg thank you so much if i could give you 1000 medals i would

OpenStudy (mathstudent55):

You're welcome.